Use elimination to solve the system of equations.
4x-9y=15
3x-4y=14

Answer by aswathytony(47)

4x - 9y = 15 ...........(1)
3x - 4y = 14 ..........(2)
elimination by equating coefficient method, make coefficient of any of the variable same & equating for the other.
i.e coefficient of x made equal by 3 x (1) & 4 x (2)
(1) x 3
4x * 3 - 9y * 3 = 15 *3
12x - 27 y = 45 ...............(3)
(2)x 4
3x * 4 - 4y *4 = 14* 4
12x - 16y = 56 ................(4)
in eq: (3) & (4) coefficient of x are same , so eliminate it by subtracting (3) & (4)
12x -27y = 45 ........(3)
12x - 16y = 56 ...........(4)
(3) -(4)
0 - 11 y = -11
-11y = -11
y = -11 /- 11
= 1.
put y=1 in (1) & find x
4x -9y = 15
4x - 9 *1 = 15
4x -9 = 15
4x= 15+9 = 24
x = 24/ 4 = 6 .